AIDA
GELINA
BRIKEN
nToF
CRIB
ISOLDE
CIRCE
nTOFCapture
DESPEC
DTAS
EDI_PSA
179Ta
CARME
StellarModelling
DCF
K40
CARME
Draft saved at 00:00:00
Fields marked with
*
are required
Entry time:
Fri May 12 16:24:20 2023
Author
*
:
Subject
*
:
<p>Since yesterday evening, the bottom right detector was at -150 V. In the morning, the current was quite stable ~ 6.77 uA. The bias was switched off, the foils opened, both of the detector arms were moved in and out three times, and the chamber was tightly covered again. A negative bias voltage was applied from 0 to 150 V in steps of 10 V one by one to all four detectors, and the V-I results can be seen in Attachments 1-4. <strong>All detectors seem to be fine.</strong> The variation of the current with time at a constant bias voltage of -150 V was checked for all detectors and the results are in Attachments 5-8. Also, with an applied bias voltage of -150 V to the bottom right and left detector, the detector arms were moved in and out three times which showed a variation of 1-2 nA and 1 nA in the leakage current. For further conditioning of the bottom right detector, it is again left with a negative bias voltage of 150 V over the weekend.</p>
Encoding
:
HTML
ELCode
plain
Suppress Email notification
Resubmit as new entry
Attachment 1:
detector_1.png
Original size: 1918x998
Attachment 2:
detector_2.png
Original size: 1918x998
Attachment 3:
detector_3.png
Original size: 1918x998
Attachment 4:
detector_4.png
Original size: 1918x998
Attachment 5:
detector_1_variation.png
Original size: 1918x998
Attachment 6:
detector_2_variation.png
Original size: 1918x998
Attachment 7:
detector_3_variation.png
Original size: 1918x998
Attachment 8:
detector_4_variation.png
Original size: 1918x998
Attachment 9:
Drop attachments here...
Draft saved at 00:00:00
ELOG V3.1.4-unknown